# DemoRank: Selecting Effective Demonstrations for Large Language Models in Ranking Task

### v1

@@ 5.2 Main Results
- w/o DReranker 51.69 68.44 44.40 54.84
- w/o DTS 52.09 67.12 46.64 55.28
DemoRank 52.52 68.28 46.90 55.90
@@
E5 49.60 66.40 39.71 51.90
$\text{DemoRank}_{\text{E5}}$ 50.74 67.37 41.76 53.29
Table 2: Results (NDCG@10) of different variants.
5.3 Analysis
In this section, we discuss different variants of DemoRank, compare DemoRank with supervised models, evaluate its performance on different demonstration numbers, and generalization on unseen datasets.
5.3.1 Different Variants of DemoRank
To understand the effectiveness of each component in DemoRank, we further evaluate different variants of DemoRank. We conduct the experiments on DL19, NQ sssand FEVER with 3-shot ICL, shown in Table 2. First, we remove our demonstration reranker DRanker and only use demonstrations retrieved by our demonstration retriever DRetriever, denoted as “- w/o DRanker”. We can see that removing DRanker causes about 1 point drop, which indicates that the reranked demonstrations are more useful for ICL. Secondly, to further validate the effectiveness of our dependency-aware training samples DTS in few-shot ICL, we introduce another variant that score each retrieved demonstration in $Z^{r}$ independently based on LLM, denoted as “- w/o DTS”. Without considering the demonstration dependency, this variant lags behind DemoRank by 0.62 points, which proves that the dependency-aware training samples align more with the few-shot ICL. Thirdly, we also replace our trained DRetriever with E5 in our framework to validate the training effectiveness of our DReranker on different demonstration retrievers, denoted as $\text{DemoRank}_{\text{E5}}$. From the results, we can see that $\text{DemoRank}_{\text{E5}}$ significantly improves E5, which proves that our DReranker’s training is flexible and not restricted by specific demonstration retriever. In addition, we also discuss the effectiveness of the ranking loss $L_{r}$ and in-batch negatives during DRetriever’s training in Appendix D.
@@ 5.3.1 Different Variants of DemoRank
DemoRank 48.14 16.90 39.76 46.47 35.93 83.96 36.14 43.90
Table 3: Results (NDCG@10) on BEIR. Best results are marked in bold. We use MS MARCO’s demonstration pool for retrieval and 3-shot ICL for E5 and DemoRank.
QNum Method MS MARCO DL19 DL20
0 0-shot 33.24 66.13 65.57
500K monoBERT 39.97 70.72 67.28
monoT5 40.05 70.58 67.33
DemoRank 35.12 68.28 67.66
20K monoBERT 30.69 63.61 59.32
monoT5 29.79 61.16 52.72
DemoRank 34.63 67.25 66.67
Table 4: Results (NDCG@10) on MS MARCO, DL19 and DL20. QNum represents the number of queries used in the MS MARCO training set.
5.3.2 Comparison with Supervised Reranker
The training of DemoRank is primarily based on queries in the training set, which can also be used to finetune a supervised model. In this part, we compare DemoRank with two supervised passage ranking models (monoBERT Nogueira and Cho (2019) and monoT5 Nogueira et al. (2020)) under different quantities of training queries. Training details of monoBERT and monoT5 are provided in Appendix B. We choose MS MARCO as the training set and NDCG@10 as the metric. We also report the 0-shot performance as a reference. The results are shown in Table 4. We can see that when provided with 500K queries, although DemoRank slightly outperforms monoBERT and monoT5 on DL20, it still lags behind them on DL19 and MS MARCO, indicating the advantages of supervised models when abundant training data is available. However, when the number of queries is limited to 20K, DemoRank significantly outperforms the two supervised models on three datasets and also shows a significant improvement over 0-shot baseline. This suggests that when training data is limited, DemoRank is more effective than supervised models, highlighting the potential of DemoRank in low-resource scenarios.
5.3.3 Different Demonstration Numbers
Demonstration number is often considered a key factor affecting ICL. In this part, we discuss the performance of our models under different demonstration numbers. We compare DemoRank with E5 baseline on FEVER and NQ datasets, using NDCG@10 as the metric. we also compare with our DRetriever to better understand the performance of our DReranker. The results are shown in Figure 3. We can see that both DRetriever and DemoRank outperform E5 consistently across different demonstration numbers, proving the effectiveness and robustness of our models. Besides, we can observe that as the demonstration number increases, the gap between DemoRank and DRetriever becomes more obvious (especially on FEVER), proving the effectiveness of dependency-aware demonstration reranking in few-shot ICL.
Figure 3: The impact of demonstration number.
5.3.4 Generalization on Unseen Datasets
@@
One of the application scenarios of DemoRank is its generalization on unseen datasets. To prove this, we evaluate DemoRank trained on MS MARCO dataset on a series of BEIR datasets. We choose 0-shot, E5 demonstration retriever, and a supervised passage ranker MonoBERT Nogueira and Cho (2019), which is also trained on the MS MARCO dataset, for comparison. We use the demonstration pool from MS MARCO due to the lack of training sets in most BEIR datasets. As shown in Table 3, DemoRank outperforms the second-best model E5, by an average of about 1 point, proving its generalization ability. Furthermore, we also draw an interesting observation: despite using demonstrations from MS MARCO, DemoRank improves the 0-shot baseline across all datasets, indicating the potential of cross-dataset demonstrations in ICL.
6 Conclusion
In this paper, we explore how to select demonstrations for passage ranking task and propose DemoRank. We first trains a demonstration retriever with multi-task learning based on LLM’s feedback. Then, an reasonable and efficient method is propose to construct dependency-aware training samples, serving as the training data of the demonstration reranker. Experiments on various ranking datasets prove the effectiveness of DemoRank. Further analysis shows the effectiveness of each proposed component, the advantages compared to supervised models, and generalization on BEIR, etc.
Limitations
In this paper, we introduce a novel demonstration selection framework DemoRank for passage ranking task. We acknowledge several limitations in this paper that present opportunities for future work. First, due to limited computational resources, we can not conduct experiments with larger LLMs, such as those with 30B or even 70B parameters. Second, our framework is limited to pointwise passage ranking and lacks discussion on how demonstrations can be selected in pairwise and listwise passage ranking, which can be a promising direction to explore.
Appendix A Implementation Details of DemoRank
The hyper-parameters for training the demonstration retriever and reranker are shown in Table 5. For the construction of training input, we pair each query with one relevant passage and one irrelevant passage respectively, thus generating two training inputs. The passages labeled with 1 in the training set are used as relevant passages and the irrelevant ones are sampled from the top-100 passages retrieved by BM25. The number of queries used in each dataset is listed in Table 6. The maximum length of the queries and passages is set to 100 and 64, respectively.
During inference, for each test query-passage pair, we first use our DRetriever to retrieve top-50 demonstrations and then rerank them using our DReranker. The top-ranked demonstrations are used for ICL. The prompt we used consists of the instruction, demonstrations (one or more), and test input. For zero-shot, no demonstrations are included. The instructions and demonstrations we used are listed in Table 7 and Table 8 respectively. The instructions are used only for each test query-passage pair and the LLM scoring process. The test inputs have the same format as the demonstration.
Appendix B Training Details of Supervised Models
For fair comparison with DemoRank, we construct the training data by pairing each query with one relevant passage and one irrelevant passage respectively. As for monoBERT Nogueira and Cho (2019), we start training from a bert-large-uncased model and use a binary classification loss to optimize the model. As for monoT5 Nogueira et al. (2020), we initialize the model with T5-base model and finetune the model using generative loss. The training parameters are the same as the original paper.
Retriever Model Reranker Model
Initialization e5-base-v2 DeBERTa-v3-base
Optimizer AdamW AdamW
Learning Rate 3e-5 1e-5
Batch Size 8 8
Warmup Steps 400 400
Train Epochs 2 2
$\lambda$ 0.2 -
Table 5: Hyperparameters for training the demonstration retriever and reranker model.
      Query Number
      FEVER       150K
      NQ       150K
      HotpotQA       150K
      MS MARCO       200K
Table 6: The number of training queries for our framework DemoRank.
Dataset Instruction
FEVER Given an article and a claim, predict whether the article is relevant to the claim by outputting either Yes or No. If the article is relevant to the claim, output Yes; otherwise, output No.
NQ Given a passage and a question, predict whether the passage is relevant to the question by outputting either Yes or No. If the passage is relevant to the question, output Yes; otherwise, output No.
HotpotQA Given a passage and a question, predict whether the passage is relevant to the question by outputting either Yes or No. If the passage is relevant to the question, output Yes; otherwise, output No.
TREC DL19 Given a passage and a query, predict whether the passage is relevant to the query by outputting either Yes or No. If the passage is relevant to the query, output Yes; otherwise, output No.
TREC DL20 Given a passage and a query, predict whether the passage is relevant to the query by outputting either Yes or No. If the passage is relevant to the query, output Yes; otherwise, output No.
MS MARCO Given a passage and a query, predict whether the passage is relevant to the query by outputting either Yes or No. If the passage is relevant to the query, output Yes; otherwise, output No.
Table 7: The instructions used for different datasets.
Dataset Demonstration Format
[…]
Table 8: The demonstration format used for different datasets.
Appendix C Transferability across different LLM Ranker
In previous experiments, we used the same LLM (Flan-T5-XL) as the demonstration scorer and passage ranker. It is unknown whether the passage ranker could be replaced with other LLMs in the inference stage. In this section, we evaluate DemoRank’s transferability across different LLM rankers on several datasets and compare with several baselines, including 0-shot, Random, K-means, BM25, and E5. We experiment with Flan-T5-XXL [^3] https://huggingface.co/google/flan-t5-xxl (larger model size) and Llama-3-8B-Instruct [^4] https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct (different model architecture) and the results are shown in Table 9. From the results, we can draw the following observations: (1) DemoRank outperforms all the baselines on Avg metric when using both Flan-T5-XXL and Llama-3-8B-Instruct as the LLM rankers, proving its strong transferability across different LLM rankers. (2) We observe that when using Flan-T5-XXL as LLM Ranker, DemoRank yields higher performance on FEVER, DL19, and MS MARCO (49.56, 68.74 and 35.90 respectively), compared with Flan-T5-XL (46.90, 68.28 and 35.12 respectively in Table 1). This shows DemoRank’s potential ability to improve passage ranking with larger-scale LLM rankers. (3) Comparing the overall 0-shot performance between Flan-T5-XL (see Table 1), Flan-T5-XXL and Llama-3-8B-Instruct, it is obvious that FlanT5 models perform better on average. This indicates that FlanT5 models are more suitable for passage ranking tasks, similar to findings from previous research Zhuang et al. (2023b).
HotpotQA NQ FEVER DL19 DL20 MS MARCO Avg
Initial Order 63.30 30.55 65.13 50.58 47.96 22.84 46.73
@@ Appendix C Transferability across different LLM Ranker
0-shot 56.64 47.61 37.38 66.22 64.30 34.29 51.07
Random 58.31 48.51 39.56 67.47 65.31 35.15 52.39
K-means 58.75 48.86 39.37 67.40 65.47 35.24 52.52
BM25 60.66 50.47 43.89 66.82 65.67 35.15 53.78
E5 60.74 50.14 43.86 66.45 65.44 34.84 53.58
DemoRank 62.25 51.68 49.56 68.74 65.90 35.90 55.67
@@
0-shot 55.93 36.24 27.53 58.47 55.10 28.09 43.56
Random 49.45 36.34 29.03 59.18 56.26 28.11 43.06
K-means 57.01 35.28 34.73 57.63 53.30 26.51 44.08
BM25 60.18 35.31 29.22 59.24 57.28 26.71 44.66
E5 60.09 36.76 28.64 57.36 53.09 26.98 43.82
DemoRank 60.89 35.47 45.24 60.36 56.45 28.54 47.83
Table 9: Results (NDCG@10) of different LLM ranker.
Appendix D Discussion on DRetriever’s Training
In this part, we conducted experiments to verify the rationale of Dretriever’s training. Firstly, we remove the ranking loss $L_{r}$ (Equation (6)) from training (denoted as “- w/o $L_{r}$”) and find a significant performance degradation on the three datasets. This indicates that the ranking signal in demonstration candidates is useful for demonstration retriever training. Besides, as we mentioned in Section 4.2, we do not apply in-batch negatives when calculating the contrastive loss $L_{c}$, which is different from previous studies Wang et al. (2023); Li et al. (2023); Karpukhin et al. (2020). To verify its rationale, we incorporate the in-batch negatives into the calculation of contrastive loss, denoted as “- w/ IBN”. From the results, we can see that the in-batch negatives do not bring significant improvement and even harm the retriever’s performance on DL19. This is because the utility of demonstrations in ranking tasks is not directly related to their similarity with the training input and the randomly sampled in-batch demonstrations may still contain valuable information and act as positive candidates, which is different from the assumption in passage retrieval Karpukhin et al. (2020). Thus, directly using in-batch negatives may introduce additional noise into the training process.
Method NQ DL19 FEVER Avg
- w/o $L_{r}$ 50.60 67.65 43.65 53.97
- w/ IBN 51.68 67.14 44.43 54.42
DRetriever 51.69 68.44 44.40 54.84
Table 10: The results (NDCG@10) of different training variants of DRetriever. We apply 3-shot ICL for each model.

### v2

@@ 5.3.1. Different Variants of DemoRank
- DRetriever w/ IBN 44.43 51.68 67.14 54.42
- DRetriever w/o $L_{r}$ 43.65 50.65 67.65 53.98
- DemoRank w/o DReranker 44.40 51.69 67.70 54.60
- DemoRank w/o Dependency 46.64 51.72 67.12 55.16
DemoRank 48.43 52.31 67.76 56.17
@@
E5 39.71 49.60 66.40 51.90
$\text{DemoRank}_{\text{E5}}$ 47.29 51.76 68.19 55.75
Table 3. Results (NDCG@10) on MS MARCO, DL19 and DL20. QNum represents the number of queries used in the MS MARCO training set.
QNum Method MS MARCO DL19 DL20
0 0-shot 33.24 66.13 65.57
500K monoBERT 39.97 70.70 67.28
monoT5 40.05 70.58 67.33
monoFLAN-T5 36.22 70.72 66.26
DemoRank 35.31 67.76 66.76
20K monoBERT 30.69 63.61 59.32
monoT5 29.79 61.16 52.72
monoFLAN-T5 32.90 64.23 64.90
DemoRank 34.97 67.33 67.53
5.3.2. Comparison with Supervised Rankers
The training of DemoRank is primarily based on queries in the training set, which can also be used to directly finetune a supervised passage ranker. MonoBERT (Nogueira and Cho, 2019) and monoT5 (Nogueira et al., 2020) are two supervised baselines widely used for comparison in passage ranking task. In this part, we choose monoBERT (340M) and monoT5 (220M) for comparison. Besides, we also introduce an LLM-based supervised ranker, named monoFLAN-T5, which uses the FLAN-T5-XL as the foundation model to enable a more fair comparison with DemoRank in LLM ranker size and adopts similar training strategy as monoT5. We compare DemoRank with the three baselines under different quantities of training queries (500K and 20K). We choose MS MARCO as the training set and NDCG@10 as the metric. We also report the 0-shot performance as a reference. The results are shown in Table 3. We can see that supervised models have a performance advantage when provided with 500K training queries. However, when the number of training queries is limited to 20K, DemoRank significantly outperforms them on three datasets and also shows a significant improvement over the 0-shot baseline. This suggests that when training data is limited, DemoRank is more effective than supervised models, highlighting the potential of DemoRank in low-resource scenarios.
Figure 3. The impact of demonstration number.
5.3.3. Different Demonstration Numbers
Demonstration number is often considered a key factor affecting ICL. In this part, we discuss the performance of our models under different demonstration numbers. We compare DemoRank with E5 baseline on FEVER and HotpotQA datasets, using NDCG@10 as the metric. we also compare with our DRetriever to better show the improvement of our DReranker. The results are shown in Figure 3. We can see that both DRetriever and DemoRank outperform E5 consistently across different demonstration numbers, and our DReranker consistently improves DRetriever, which proves the effectiveness of our proposed models. Besides, we observe that when expanding from 2-shot to 3-shot, the performance of DemoRank on HotpotQA slightly decreases. We speculate that this is because 2-shot demonstrations already provide sufficient knowledge for the model, and adding more demonstrations may introduce unnecessary noise.
Table 4. Results (NDCG@10) on BEIR. Best results are marked in bold. We use MS MARCO’s demonstration pool for retrieval and 3-shot ICL for E5 and DemoRank.
@@ 5.3.3. Different Demonstration Numbers
DemoRank 48.94 17.09 39.67 46.66 36.18 84.04 36.05 44.09
5.3.4. Generalization on Unseen Datasets
@@
One of the application scenarios of DemoRank is its generalization on unseen datasets. To prove this, we evaluate DemoRank trained on MS MARCO dataset on a series of BEIR datasets. We choose 0-shot, E5 demonstration retriever, and the supervised passage ranker MonoBERT (Nogueira and Cho, 2019), which is also trained on the MS MARCO dataset, for comparison. We use the demonstration pool from MS MARCO due to the lack of training sets in most BEIR datasets. The results are shown in Table 4. We can see that DemoRank outperforms the second-best model E5, by an average of about 1 point, proving its generalization ability. Furthermore, we also draw an interesting observation: despite using demonstrations from MS MARCO, DemoRank improves the 0-shot baseline across all datasets, indicating the potential of cross-dataset demonstrations in ICL.
Table 5. Results (NDCG@10) of different LLM ranker. We apply 3-shot ICL for BM25, E5 and DemoRank.
Method FEVER NQ DL19 Average
Initial Order 65.13 30.55 50.58 46.73
@@
0-shot 37.38 47.61 66.22 50.40
BM25 43.89 50.47 66.82 53.73
E5 43.86 50.14 66.45 53.48
DemoRank 50.89 51.68 69.23 57.27
@@
0-shot 27.53 36.24 58.47 40.75
BM25 29.22 35.31 59.24 41.26
E5 28.64 36.76 57.36 40.92
DemoRank 45.50 37.11 60.93 47.85
5.3.5. Transferability across different LLM Rankers
In previous experiments, we used the same LLM (Flan-T5-XL) as the demonstration scorer and passage ranker. It is unknown whether the passage ranker could be replaced with other LLMs in the inference stage. In this section, we evaluate DemoRank’s transferability across different LLM rankers on FEVER, NQ and DL19 datasets, and compare with several baselines including 0-shot, BM25, and E5. We experiment with Flan-T5-XXL (larger model size) and Llama-3-8B-Instruct (different model architecture) and the results are shown in Table 5. From the results, we can see that DemoRank outperforms all the baselines when using two different LLM rankers, proving its strong transferability. Besides, we observe that when using Flan-T5-XXL as LLM Ranker, DemoRank yields higher performance on FEVER and DL19 (50.89 and 69.23) compared with Flan-T5-XL (48.43 and 67.76 in Table 1). This shows DemoRank’s potential ability to improve passage ranking with larger-scale LLM rankers.
Figure 4. The impact of retrieved demonstration number on the performance of DemoRank.
5.3.6. The impact of retrieved demonstration number $D$
Our DReranker aims to select few-shot demonstrations from top-$D$ demonstrations retrieved by DRetriever. Thus, DReranker’s performance highly relies on the quality of these retrieved $D$ demonstrations. In this part, we intend to explore the impact of different retrieval numbers $D$ on the performance of DReranker. We choose different retrieval numbers $D$ (10, 20, 30 and 40), perform reranking using DReranker, and evaluate the performance of 3-shot ICL on FEVER and HotpotQA datasets. We also report the results of DRetriever as a reference. The results are shown in Figure 4. From the results, we observe that increasing the retrieval number $D$ consistently enhances the DemoRank’s performance on both datasets. This indicates that a larger pool of retrieved demonstrations allows the DReranker to select more effective few-shot demonstrations. Besides, DemoRank’s performance on the HotpotQA dataset shows marginal improvement when the retrieval number increases from 30 to 40. This suggests that the most useful demonstrations are primarily concentrated within the top 30 retrieved ones on HotpotQA.
6. Conclusion
In this paper, we discuss the challenges of demonstration selection in ranking task and propose DemoRank framework which consists of a demonstration retriever and a dependency-aware reranker. Different from the demonstration retriever’s training which models demonstrations independently, the demonstration reranker is trained using dependency-aware training samples which are constructed using an efficient method. We also design a novel list-pairwise training approach for reranker optimization, which compares a pair of demonstration lists that differ only in the last demonstration. Experiments on various ranking datasets prove the effectiveness of DemoRank. Further analysis shows the effectiveness of each proposed component, the advantages compared to supervised models, performance on different demonstration numbers, generalization on unseen datasets, etc.